1. A discrete random variable 𝑋 has the following probability mass function

𝑃(𝑋=π‘₯)={2π‘˜π‘₯ π‘₯=2,4,6

{π‘˜(π‘₯+2) π‘₯=8

{ 0 π‘œπ‘‘β„Žπ‘’π‘Ÿπ‘€π‘–π‘ π‘’

where π‘˜ is a constant

a. Show that π‘˜=1/34

b. Find the exact value of 𝑃(4<π‘₯≀8)

c. Find the exact value of 𝑃(2<π‘₯<4)

d. What is the expected value of the random variable 𝑋?

e. What is the variance of the random variable 𝑋?

f. Determine π‘‰π‘Žπ‘Ÿ(5βˆ’3𝑋)


1
Expert's answer
2021-04-27T17:07:07-0400

"P(X = x) = \left\{ {\begin{matrix}
{2kx,\,\,x = 2,4,6}\\
{k(x + 2),\,\,x = 8}\\
{0,\,\,otherwise}
\end{matrix}} \right."

Find the values ​​of the probabilities

"P(x = 2) = 2k \cdot 2 = 4k"

"P(x = 4) = 2k \cdot 4 = 8k"

"P(x = 6) = 2k \cdot 6 = 12k"

"P(x = 8) = k \cdot (8 + 2) = 10k"

a) We find the value of k from the condition

"\sum {{P_i}} = 1"

Then

"4k + 8k + 12k + 10k = 1 \Rightarrow 34k = 1 \Rightarrow k = \frac{1}{{34}}"

Q.E.D

b) construct a series of probability distributions

"\begin{matrix}
X&2&4&6&8\\
P&{\frac{4}{{34}}}&{\frac{8}{{34}}}&{\frac{{12}}{{34}}}&{\frac{{10}}{{34}}}
\end{matrix}"

Let's find

"P(4 < x \le 8) = P(6) + P(8) = \frac{{12 + 10}}{{34}} = \frac{{11}}{{17}}"

Answer: "P(4 < x \le 8) = \frac{{11}}{{17}}"

c) Let's find

"P(2 < X < 4) = 0" (since, by condition, "P(X = x) = 0" for all "x \in \left( {2;4} \right)" )

Answer: "P(2 < X < 4) = 0"

d) Let's find he expected value:

"M(x) = \sum {{x_i}} {p_i} = \frac{{2 \cdot 4 + 4 \cdot 8 + 6 \cdot 12 + 8 \cdot 10}}{{34}} = \frac{{96}}{{17}}"

Answer: "M(x) = \frac{{96}}{{17}}"

e) Let's find the variance:

"Var(x) = M({x^2}) - {M^2}(x) = \frac{{4 \cdot 4 + 16 \cdot 8 + 36 \cdot 12 + 64 \cdot 10}}{{34}} - {\left( {\frac{{96}}{{17}}} \right)^2} = \frac{{1120}}{{289}}"

Answer: "Var(x) = \frac{{1120}}{{289}}"

f) Let's find

"Var(5 - 3x) = Var(5) + Var( - 3x) = 0 + {( - 3)^2}Var(x) = 9 \cdot \frac{{1120}}{{289}} = \frac{{{\rm{10080}}}}{{289}}"